1 litre of gas at STP is expanded adibatically to 3 litre.Find work done by gas.gaama=1.4&3power1.4=4.65

Grade:12

1 Answers

Arun
3 years ago
V1=1l initial volume, V1=1×103m3
V2=3l final volume, V2=3×103m3
At STP:
T=273K
P1=101.325kPa initial pressure 
Y=1.4
Work done in adiabatic process, ΔW=(Y1P2V2P1V1)
adiabatic PVγ=constant
P1V1γ=P2V2γ
P2=P1(V2V1)γ=1.01325(3×1031×103)1.4
 
P2=101.325×(3)1.41=(4.6555101.325)kPa
 
P2=21.7646KPa
 
ΔW=(1.4121.7646×103×3×103101.325×103×1×103)
 
ΔW=(0.465.2937101.325)=90.5J
Considering only magnitude, the work done by the air will be 90.5J